HUBBLE SPACE TELESCOPE DAILY REPORT #1227 PERIOD COVERED: 0000Z Daily Status Report as of 263/0000Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1C: 4-Chip STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed FOS/BL 5610 (Symbiotic Systems in the Magellanic Clouds Cycle 4 High) The Faint Object Spectrograph (Blue Side) was used to make ultraviolet observations of the symbiotic system S63 located in one of the Magellanic clouds to learn more about the hot components in this system, such as radiation temperature and the luminosity. The observations were completed as planned, and no problems were reported. 1.3 Completed GHRS/2 5414 (Dayglow Emission Line Profile from the Outer Planets: cycle 4 Medium) The High Resolution Spectrograph (Side Two) was used to make observations of the Lyman-Alpha profiles of the moon Triton on Neptune. A wavelength calibration observation was also taken. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 5571 (WF/PC-2 Cycle 4 Calibration: Earthflats, Smaller Set) The WF/PC-2 was used to make several ultraviolet and narrow-band filter Earth flat calibration observations to correct for the telescope's illumination pattern. The observations were completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 5493 (Synoptic Monitoring of Seasonal Phenomena on Mars: Cycle 4 Medium) The WF/PC-2 was used to make observations of the planet Mars at selected seasons, looking for periodic changes. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed FGS 5554 (HIPPARCOS Plate Scale: Cycle 4 Calibration) FGS#3 was used to make astrometric observations of stars HIC22604, H39-385-RE, H39-366-RE, H39-255-RE, HIC22507-R which were observed by the HIPPARCOS satellite. The astrometric observations were monitored while telemetry was available. 1.7 Completed WF/PC-2 5370 (HST Medium-Deep Survey: Cycle 4 High, High Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations as part of a medium-deep sky survey. The observations were completed as planned, and no problems were reported. 1.8 Completed FOS/BL 5182 (UV Spectroscopy of Low Redshift Active Quasars: Cycle 4) The Faint Object Spectrograph (Blue Side) made ultraviolet observations of the low redshift active quasar AKN120 to obtain the profiles of the emission lines, and the ratios of their fluxes. HSTAR 4915 was written for this proposal since the binary search for the target star failed. 1.9 Completed WF/PC-2 5092 (GTO Low Latitude Parallel Program: Cycle 4) The WF/PC-2 was used to make observations taken in parallel with the prime science observations. This is part of a survey to take pictures of a large number of fields within 30 degrees of the galactic plane. The observations were completed as planned, and no problems were reported. 1.10 Completed FOS/RD 5588 (FOS Prism Snapshot Survey) The Faint Object Spectrograph (Red Side) was used to make a Prism observation of quasi-stellar object AKN120 to study its optical to ultraviolet continuum shape. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C135.0 as part of a continuum survey and to study the galaxy's morphology. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 5678 (Resolution of the Upper IMF in the Magellanic Clouds: Cycle 4) The WF/PC-2 was used to make observations of the bright O class star HD32228 which is located in the Magellanic Cloud to determine if this star actually contains combinations of unresolved stars. The observations were executed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled :13 Successful :13 2.3 Significant HSTARs: None 2.4 Operations Notes: There were no anomalies reported in any subsystem aboard the HST. A flight software patch was uplinked yesterday modifying the SA Post Maneuver Timeout sequence. Subsequent SA slew activity indicated that the new sequence performed nominally. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /CAW -END-